two balls A and B simutaneously thrown .A is thrown simutaneously from the ground with a velocity of 20 m/s in the upward direction and B is thrown from the height of 40 m in the downwars direction with same velocity where will the 2 balls meet?

Grade:11

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
Let height of meeting point from ground=h and time taken=t
for ball thrown upwards, h=ut-1/2 gt^2
and for ball thrown downwards, 40-h=ut+1/2 gt^2
adding both equations we get 2ut=40
or t= 40/2u=1 second
so from 1st equation h=20*1 - 1/2 * 9.8*1^2
=15.1m from ground
So the balls collide at height 15.1 m from ground. [Ans]
 
Regards,
Thank you,
Soumya Ranjan Mohanty.